OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, -3) in slope-intercept form.

OpenStudy (anonymous):

A. y = 7x - 10 B. y = 7x + 10 C. y = 7x + 18 D. y = 7x - 18

OpenStudy (anonymous):

@FireKat97

OpenStudy (anonymous):

This is the last one slope is 7

OpenStudy (firekat97):

Yupp

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] Find the slope first

OpenStudy (anonymous):

its 7

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (firekat97):

yup, thats correct, now reapply the equation from earlier

pooja195 (pooja195):

now use this

OpenStudy (anonymous):

i dont know how to get the intercept

OpenStudy (firekat97):

you don't need to find the intercept, you just have to substitute any set of co-ordinates and the slope into the (y - y1) =m(x - x1) equation

OpenStudy (firekat97):

The y intercept form is just the name of the form which follows this format y = mx + b

OpenStudy (firekat97):

So thats how you have to re-arrange the equation once you sub your slope and co-ordinates in

OpenStudy (anonymous):

I dont know what b is though

OpenStudy (anonymous):

I dont get how to figure out if its 10 or 18

OpenStudy (anonymous):

4+3=7(2-1)

OpenStudy (anonymous):

Thats what i got what i plugges it in

OpenStudy (anonymous):

plugged*

OpenStudy (anonymous):

i just get 7=7

OpenStudy (firekat97):

But you only need to plug in one set of the co-ordinates, so either (2, 4) or (-1, 3) into x1 and y1

OpenStudy (firekat97):

Just pick the set which looks to have the "easier numbers" so smaller numbers etc

OpenStudy (firekat97):

But it shouldn't matter which of the co-ordinates you choose

OpenStudy (anonymous):

okay so( x+3)=7(x-1)

OpenStudy (firekat97):

Yup, right hand side is correct, but with the left hand side, that should be a y, not an x

OpenStudy (anonymous):

whoops haha

OpenStudy (firekat97):

hahaha dw all good :)

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

haha good!

OpenStudy (firekat97):

So, now we can rearrange this equation to put it in the form of y = mx + b

OpenStudy (firekat97):

so we want to move everything except the y to the right hand side

OpenStudy (anonymous):

Its A

OpenStudy (firekat97):

yup :D

OpenStudy (anonymous):

Thank you!

OpenStudy (firekat97):

no problem :)